# RoboEXP: Action-Conditioned Scene Graph via Interactive Exploration for Robotic Manipulation

### v1

@@ III-B Interactive Exploration
This subsection describes how we can construct a complete action-conditioned scene graph of a real-world scene. This is a challenging problem due to partial observability. For instance, a banana cannot be populated without opening the cabinet. To solve this task, we formulate the scene graph construction as an active perception and exploration problem using POMDP-inspired notations. Formally, at each time $t$, based on our past graph estimation $\mathbf{G}^{t-1}$, and past sensor observations $\mathbf{O}^{t-1}$, our agent takes an action $\mathbf{A}^{t}$, which causes the environment to transition to a new state, and the agent receives a new observation $\mathbf{O}^{t}$, which is used to update its current inferred graph $\mathbf{G}^{t}$. This update might include adding new nodes to the graph or updating the state of an existing node. We will then continue with exploration and keep updating the set of remaining unexplored nodes $\mathbf{U}\subset\mathbf{V}$ (see Algorithm 1).
@@
4:      add spatial relations % memory
@@
21:         add action preconditions % memory
@@
The goal of the exploration is simple: discover and explore all the nodes of the scene graph in as little time as possible. Towards this, we formulate a reward function with three terms:
$\mathbf{R}^{t}=\mathbf{R}_{\text{graph}}^{t}+\mathbf{R}_{\text{explore}}^{t}+\mathbf{R}_{\text{time}}^{t}$
where $\mathbf{R}_{\text{graph}}^{t}=|\mathbf{V}^{t}|-|\mathbf{V}^{t-1}|$ is the graph construction term, which promotes our agent to discover as many nodes as possible to the graph, $\mathbf{R}_{\text{explore}}^{t}=\max(0,|\mathbf{U}^{t-1}|-|\mathbf{U}^{t}|)$ gives positive reward to actions that reduce unexplored node set, which prioritize the agent to explore previously unexplored nodes, and immediate reward $\mathbf{R}_{\text{time}}^{t}=-\lambda,0<\lambda<1$ is a negative time reward that optimizes the time efficiency and allows the exploration to terminate when there is no more node to explore.
IV Method
In this section, we outline the structure of our RoboEXP system, including perception, memory, decision-making, and action modules, in Sec. IV-A. We then discuss our system’s design for the interactive scene exploration task in Sec. IV-B, focusing on its application in closed-loop exploration processes that may require multi-step or recursive reasoning and handle potential interventions.
IV-A RoboEXP System
To tackle the task outlined in Section Sec. III, we present our RoboEXP system, designed to autonomously explore unknown environments by observing and interacting with them. The system comprises four key components: perception, memory, decision-making, and action modules (see Fig. 3). Raw RGBD images are captured through the wrist camera in different viewpoints and processed by the perception modules to extract scene semantics, including object labels, 2D bounding boxes, segmentations, and semantic features. The obtained semantic information is then transmitted to the memory module, where the 2D data is merged into the 3D representation. Such 3D information serves as a valuable guide for the decision module, aiding in the selection of appropriate actions to further interact or observe the environment and unveil hidden objects. The action module is activated to execute the planned action, generating new observations for the perception modules. This closed-loop system ensures the thoroughness of our task in interactive scene exploration.
Perception Module. Given multiple RGBD observations from different viewpoints, the objective of the perception module (Fig. 3a) is to detect and segment objects while extracting their semantic features. To enhance generality, we opt for the open-vocabulary detector GroundingDINO [27] and the Segment Anything in High Quality (SAM-HQ) [28], an advanced version of SAM [29]. For the extraction of semantic features used in subsequent instance merging within the memory module, we employ CLIP [30]. To obtain per-instance CLIP features, we implement a strategy similar to the one proposed by Jatavallabhula et al. [118]. Specifically, we extend the local-global image feature merging approach by incorporating additional label text features to augment the semantic CLIP feature for each instance. Furthermore, we exclusively focus on instance-level features, disregarding pixel-level features, thereby accelerating the entire semantic feature extraction process.
Memory Module. The memory module (Fig. 3b) is designed to construct our ACSG of the environment by assimilating observations over time. For the low-level memory, to ensure stable instance merging from 2D to 3D, we employ a similar instance merging strategy as presented in Lu et al. [119], consolidating observations from diverse RGBD sources across various viewpoints and time steps. In contrast to the original algorithm, which considers only 3D IoU and semantic feature similarity we additionally incorporate label similarity and instance confidence. To enhance algorithm efficiency, we represent low-level memory using a voxel-based representation, which allows for more efficient computation and memory updates. Meanwhile, given the crowded nature of objects in our tabletop setting, we have implemented voxel-based filtering designs to obtain a cleaner and more complete representation of the objects for storage in our memory.
The memory module handles merging across different viewpoints and time steps. To merge across different viewpoints, we project 2D information (RGBD, semantic features, mask, bounding box) to 3D and leverage the instance merging strategy mentioned earlier to attain consistent 3D information. Addressing memory updates across time steps presents a challenge due to dynamic changes in the environment. For instance, a closed door in the previous time step may be opened by our robot in the current time step. To accurately reflect such changes, our algorithm evaluates whether elements within our memory have become outdated, primarily through depth tests based on the most recent observations. This process ensures that the memory accurately represents the environment’s current state, effectively managing scenarios where objects may change positions or states across different time steps.
For the high-level graph of our ACSG, the memory module analyzes the relationships between objects and the logical associations between actions and objects. Depending on changes in low-level memory and relationships, the memory module is tasked with updating the graph. This involves adding, deleting, or modifying nodes and edges within our graph.
Decision-Making Module. The primary goal of the decision module (Fig. 3c) is to identify the appropriate object and corresponding skill to enhance the effectiveness and efficiency of interactive scene exploration. In the context of our task, distinct objects may necessitate distinct exploration strategies. While humans can easily discern the most suitable skill to apply (e.g., picking up the top Matryoshka doll to inspect its contents), achieving such decisions through heuristic-based methods is challenging. The utilization of a Large Multi-Modal Model (LMM), such as GPT-4V [31, 32], shows instrumental in addressing this difficulty, as it captures commonsense knowledge that facilitates decision-making.
The LMM brings commonsense knowledge to our decision-making process and serves in two pivotal roles. Firstly, it functions as an action proposer. Given the current digital environment from the memory module, GPT-4V is tasked with selecting the appropriate skill for unexplored objects in our system. For instance, when presented with a visual prompt of an object within a green bounding box from various viewpoints, GPT-4V can discern the suitable “pick up” skill for the Matryoshka doll in the environment. For unexplored objects, our ACSG includes the attribute of whether each object node is explored or unexplored. GPT-4V, in its role as the proposer, also functions to assess whether the object holds value for further exploration. If not, the corresponding node is marked as explored, indicating that no further actions are needed.
Secondly, the LMM also serves as the action verifier. For the proposer role, it analyzes the object-centric attributes and doesn’t consider surrounding information when choosing the proper skill. For example, if the proposed action involves opening a door, the proposer alone may struggle with cases where obstructions exist in front of the door (e.g., a condiment bottle). To address this, we use another LMM program to verify the feasibility of the action and identify any objects in the scene that may impede the action based on information from our ACSG.
In summary, the decision module, with its dual roles, effectively guides our system to choose efficient actions that minimize uncertainty in the environment and successfully locate all relevant objects.
Action Module. In the action module (Fig. 3d), our primary focus is on autonomously constructing the ACSG through effective and efficient interaction with the environment. We employ heuristic-based action primitives within our action module, leveraging the geometry cues in our ACSG. These primitives encompass seven categories: “open the door”, “open the drawer”, “close the door”, “close the drawer”, “pick object to idle space”, “pick back object”, “move wrist camera to position”. Strategic utilization of these skills plays a pivotal role in accomplishing intricate tasks seamlessly within our system (more details in the Appendix).
IV-B Other Design in Interactive Exploration
One desiderata for robot exploration is the ability to handle scenarios that necessitate multi-step or recursive reasoning. An example of this is the Matryoshka doll case (Fig. 6b), which cannot be addressed using previous one-step LLM-based code generation approaches [34, 33]. In contrast, our modular design allows agents to dynamically plan and adapt in a closed-loop manner, enabling continuous LLM-based exploration based on environmental feedback.
V Experiments
In this section, we assess the performance of our system across a variety of tabletop scenarios in the interactive scene exploration setting. Our primary objective is to address two key questions through experiments: 1) How does our system effectively and efficiently deal with diverse exploration scenarios and successfully construct comprehensive ACSG? 2) What is the utility of our ACSG in facilitating downstream tasks?
V-A Robot and Environment Setups
All our experiments are conducted in a real-world setting. In these scenarios, we mount one RealSense-D455 camera on the wrist of the robot arm to collect RGBD observations, with the execution of actions performed by the UFACTORY xArm 7. The end effector for our robot arm is the soft gripper (see Fig. 3). Our experimental setup encompasses a diverse range of objects, as illustrated in Fig. 4.
V-B Interactive Exploration and Scene Graph Building
To assess our system’s efficacy across various exploration scenarios, we compared it with a strong baseline by augmenting GPT-4V with ground truth actions. We designed five types of experiments, each with 10 different settings varying in object number, type, and layout (see the Appendix for more details). Our quantitative analysis reveals that our RoboEXP system consistently surpasses the baseline across various tasks. Furthermore, we validate the performance of our system in constructing ACSG through qualitative demonstrations.
Baseline. We employ the pure GPT-4V as our baseline model along with the chain-of-thoughts (CoT) to enhance its capabilities, as outlined in a method similar to that proposed by Hu et al. [34]. This baseline operates in a closed-loop fashion, receiving three RGB observations from different viewpoints during each iteration. At each turn, it generates the current scene graph, encompassing hidden objects, and suggests the next action to be taken. Upon determining that all tasks are completed, the model outputs “Done” (refer to the complete prompts in the Appendix). To ensure the baseline is robust, we utilize manual actions as ground truth references for the proposed actions. For instance, if the baseline suggests opening a specific drawer, we manually perform the action and prompt the model with the new observation to generate another action. In contrast, in the exploration experiments described below, all actions from our system are automatically executed by our action module on the physical robot.
Evaluation. To thoroughly assess the efficacy of our system compared to the baseline, we have designed five key metrics to measure its performance. It is crucial to note that the output of our task, represented by ACSG, aligns precisely with the format of ACSG for our system. Conversely, for the baseline, we manually construct ACSG based on its actions and the new observations it uncovers. Due to the unstructured nature of the raw scene graph from the baseline, we carefully refine it according to the observable objects, providing an upper-bound baseline for comparison during evaluation.
To assess the effectiveness and efficiency of ACSG, we engage human evaluators in the tasks to construct the ground truth version of ACSG. The five main metrics employed for evaluation are as follows:
These metrics provide a comprehensive evaluation of the method’s performance. Additionally, we visualize the number of objects and actions during the exploration process to show the exploration strategies employed by different methods.
TABLE I: Quantitative Results on Different Tasks. We compare the performance of both the GPT-4V baseline and our system across various tasks. We assess the outcomes using five distinct metrics to illustrate diverse facets of the interactive exploration process. Our system consistently outperforms the baseline across all tasks and metrics.
Metric GPT-4V Ours GPT-4V Ours GPT-4V Ours GPT-4V Ours GPT-4V Ours
[…]
Figure 5: Visualization of Quantitative Results. (a) The action-object graph captures the change in the number of discovered objects relative to the number of actions taken. Our RoboEXP efficiently discovers all objects. Sometimes, the object count doesn’t increase during actions due to the absence of objects in storage after opening. Additionally, some actions are employed to restore the scene state (e.g., closing the door after exploration). (b) The error breakdown of all our quantitative experiments includes 5 task settings with 10 variations each. We categorize errors into perception, decision, action, and no-error cases. For the GPT-4V baseline, manual assistance in action execution eliminates failure cases, serving as an upper bound for baseline performance. Even in this scenario, our RoboEXP largely outperforms the baseline.
Comparison. The quantitative findings presented in Tab. I underscore the superior performance of our system compared to the baseline method. Our approach showcases a notable enhancement across all metrics, outperforming the baseline by a considerable margin. The collective assessment of success rate, object recovery, and unexplored space metrics unequivocally validates the efficacy of our system in exploring unfamiliar scenes through interactive processes. It is essential to highlight that in the case of object recovery, the baseline method may occasionally choose to randomly open certain drawers or doors to unveil objects. This randomness contributes to a seemingly higher object recovery rate for the baseline, which may not necessarily correlate with its overall success. The unexplored space metric shows that our system is much more stable in exploring all need-to-explore spaces.
Moreover, both the success rate and graph edit distance underscore the close alignment of our system with human actions, highlighting the efficiency of our approach across diverse scenarios. The state recovery metric assesses whether the final state post-exploration resembles the initial state. Our system consistently shows effective state recovery; however, the baseline may trick this metric by opting not to take any action, resulting in an artificially high score in this aspect.
V-C Utility of our ACSG
The scenarios depicted in Fig. 1 exemplify the efficacy of our generated output (ACSG) in manipulation tasks. Consider the table-rearranging scenario: without our ACSG, the robot struggles to swiftly prepare the table due to the lack of precise prior knowledge about the location of objects (e.g., the fork stored in the top-left drawer of the wooden cabinet). Beyond comprehensive layout guidance, our ACSG also addresses a crucial question regarding task feasibility for the robot. For instance, if there is no spoon in the scene, the robot recognizes its inability to perform the task and asks for human help.
In addition to enhancing downstream manipulation tasks, our ACSG possesses the capability to autonomously adapt to environmental changes. In the human intervention setting, our system seamlessly explores newly added components, such as a cabinet, ensuring continuous adaptability. Check our Appendix and supplemental video for more details.
V-D Remaining Challenges
Although our system has proven effective, there is room for improvement. The breakdown of the failure rate in Fig. 5.b suggests that failures primarily arise from detection and segmentation errors within the perception module. To address this issue, we envision two future directions: 1) enhancing the capabilities of visual foundation models for open-world semantic understanding, and 2) utilizing temporal cues and semantic fusion techniques to improve perception robustness through continuous observations.
Furthermore, our system would benefit from enhanced LMM capacities and the integration of sophisticated skill modules, including learning-based or model-based path planning. Such improvements would improve both the decision-making and action modules, thereby further reducing failure cases.
VI Conclusion
We introduced RoboEXP, a foundation-model-driven robotic exploration framework capable of effectively identifying all objects in a complex scene, both directly observable and those revealed through interaction. Central to our system is action-conditioned 3D scene graph, an advanced 3D scene graph that goes beyond traditional models by explicitly modeling interactive relations between objects. Experiments have shown RoboEXP’s superior performance in interactive scene exploration across various challenging scenarios, significantly outperforming a strong GPT4V-based baseline. Notably, the reconstructed action-conditioned 3D scene graph is crucial for guiding complex downstream manipulation tasks, like preparing breakfast in a mock-kitchen environment with fridges, cabinets, and drawer sets. Our system and its action-conditioned scene graph lay the groundwork for practical robotic deployment in complex settings, especially in environments like households and offices, facilitating their everyday use.
Acknowledgements
Yunzhu Li is partly supported by the Amazon AICE Award. Shenlong Wang is supported by the Amazon AICE Award, IBM IIDAI Grant, Nvidia Hardware Grants, the Insper-Illinois Innovative Grant, the NCSA Faculty Fellow, NSF Award #2331878 and #2312102, and gifts from Intel and Meta. This work does not relate to the positions of Shubham Garg and Hooshang Nayyeri at Amazon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
Appendix
I Additional Details of RoboEXP system
I-A Interactive Exploration
Due to space constraints, we did not include a comprehensive explanation of the algorithm proposed in the problem statement, but include more details here for clarity. We formulate the interactive scene exploration task into an active perception and exploration problem to construct the action-conditioned 3D scene graph (ACSG).
@@ I-A Interactive Exploration
The algorithm shown in the main paper simply mentions “add spatial relations” and “add action preconditions” as part of the function of the memory module, but without detailed explanation. In the algorithm, we have demonstrated how to construct the edges from objects to actions $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{a}}$ and from actions to objects $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{a}}$; however, there is a lack of description for the other two types of edges.
@@
Add Spatial Relations. The logic involves analyzing the spatial relationships among objects using spatial heuristics and incorporating the resulting spatial relation edges between objects $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{o}}$ (see Algorithm 2).
@@
Add Action Preconditions. The approach is to assess the feasibility of implementing the actions. We utilize the decision-making module to verify whether there are any prerequisite actions that need to be completed beforehand, and then adjust the plan accordingly (see Algorithm 3).

### v2

@@ 3 Action-Conditioned 3D Scene Graph
Interactive Exploration. Constructing a complete ACSG of a real-world scene is challenging due to partial observability. To solve this task, we formulate the scene graph construction as an active perception and exploration problem using POMDP-inspired notations. Formally, at each time $t$, based on our past graph estimation $\mathbf{G}^{t-1}$, and past sensor observations $\mathbf{O}^{t-1}$, our agent takes an action $\mathbf{A}^{t}$, which causes the environment to transition to a new state, and the agent receives a new observation $\mathbf{O}^{t}$, which is used to update its current inferred graph $\mathbf{G}^{t}$. This update might include adding new nodes to the graph or updating the state of an existing node. We will then continue with exploration and keep updating the set of remaining unexplored nodes $\mathbf{U}\subset\mathbf{V}$ (See the Appendix for the full algorithm). The goal of the exploration is simple: discover and explore all the nodes of the scene graph in as little time as possible. We thus formulate a reward function with three terms: $\mathbf{R}^{t}=\mathbf{R}_{\text{graph}}^{t}+\mathbf{R}_{\text{explore}}^{t}+\mathbf{R}_{\text{time}}^{t}$, where $\mathbf{R}_{\text{graph}}^{t}=|\mathbf{V}^{t}|-|\mathbf{V}^{t-1}|$ promotes our agent to discover as many nodes as possible to the graph, $\mathbf{R}_{\text{explore}}^{t}=\max(0,|\mathbf{U}^{t-1}|-|\mathbf{U}^{t}|)$ gives a positive reward to actions that reduce the unexplored node set, which prioritizes the agent to explore previously unexplored nodes (note that an action can lead to the discovery of new unexplored nodes being added to the graph), and immediate reward $\mathbf{R}_{\text{time}}^{t}=-\lambda,0<\lambda<1$ is a negative time reward that enhances efficiency and terminate the exploration when there is no more node to explore.
4 Interactive Robotic Exploration (RoboEXP) System
In this section, we outline the structure of our RoboEXP system, including perception, memory, decision-making, and action modules (See the Appendix for full details of our implementation).
Figure 3: Overview of Our RoboEXP System. We present a comprehensive overview of our RoboEXP system, comprised of four modules: (a) perception, (b) memory, (c) decision-making and (d) action module.
Perception Module. Given multiple RGBD observations from different viewpoints, the objective of the perception module (Fig. 3a) is to detect and segment objects while extracting their semantic features. To enhance generality, we opt for the open-vocabulary detector GroundingDINO [91] and the Segment Anything in High Quality (SAM-HQ) [92], with a predefined list of tags. For the extraction of semantic features used in subsequent instance merging, we employ per-instance CLIP [93], following a similar strategy to the one proposed by Jatavallabhula et al. [94].
Memory Module. The memory module (Fig. 3b) handles merging across different viewpoints and time steps. To merge across different viewpoints, we project 2D information to 3D and leverage the instance merging strategy similar to Lu et al. [95] to attain consistent 3D information. Addressing memory updates across time steps presents a challenge due to dynamic changes in the environment. For instance, a closed door in the previous time step may be opened by our robot in the current time step. To accurately reflect such changes, our algorithm evaluates whether elements within our memory have become outdated, primarily through depth tests based on the most recent observations. This process ensures that the low-level memory accurately represents the environment’s current state, effectively managing scenarios where objects may change positions or states across different time steps. For the high-level graph of our ACSG, the memory module analyzes the relationships between objects and the logical associations between actions and objects. Depending on changes in low-level memory and relationships, the memory module is tasked with updating the graph. This involves adding, deleting, or modifying nodes and edges within our graph.
Decision-Making Module. The primary goal of the decision module (Fig. 3c) is to identify the appropriate object and corresponding skill to enhance the effectiveness and efficiency of interactive scene exploration. In the context of our task, distinct objects may necessitate distinct exploration strategies. While humans can easily discern the most suitable skill to apply (e.g., picking up the top Matryoshka doll to inspect its contents), achieving such decisions through heuristic-based methods is challenging. The LMM brings commonsense knowledge to our decision-making process and serves in two pivotal roles. Firstly, it functions as an action proposer. Given the current digital environment from the memory module, GPT-4V is tasked with selecting the appropriate skill for unexplored objects in our system. Secondly, the LMM also serves as the action verifier. For the proposer role, it analyzes the object-centric attributes but ignore surrounding information when choosing the proper skill. To address this, we use another LMM program to verify the feasibility of the action and identify any objects in the scene that may impede the action based on information from our ACSG.
Action Module. In the action module (Fig. 3d), our primary focus is on autonomously constructing the ACSG through effective and efficient interaction. We design a set of action primitives that operate on the low-level geometric information embedded within our ACSG. These primitives encompass seven categories: “open the door”, “open the drawer”, “close the door”, “close the drawer”, “pick object to idle space”, “pick back object”, “move wrist camera to position”. Strategic utilization of these skills plays a pivotal role in accomplishing intricate tasks seamlessly within our system.
5 Experiments
In this section, we assess the performance of our system across a variety of tabletop scenarios. Our primary objective is to address two key questions through experiments: 1) How does our proposed system handle diverse exploration scenarios well and build the complete ACSG? 2) How useful is our ACSG in the downstream object retrieval and manipulation tasks?
5.1 Interactive Exploration and Scene Graph Building
Robot Setups. All our experiments are conducted in a real-world setting. In the tabletop scenarios, we mount one RealSense-D455 camera on the wrist of the robot arm to collect RGBD observations, with the execution of actions performed by the UFACTORY xArm 7. The end effector for our robot arm is the soft gripper (see Fig. 3). In the mobile setting, we choose the Hello Robot Stretch2 with the official upgraded kits. See the Appendix for more details on the full setup.
Table 1: Quantitative Results on Different Tasks. We compare the performance of both the GPT-4V baseline and our system across various tasks. Our system consistently outperforms the baseline across all metrics.
Metric GPT-4V Ours GPT-4V Ours GPT-4V Ours GPT-4V Ours GPT-4V Ours
[…]
Figure 4: Visualization of Quantitative Results. (a) The action-object graph captures the change in the number of discovered objects relative to the number of actions taken. Our RoboEXP efficiently discovers all objects. (b) The error breakdown of all our quantitative experiments includes 5 task settings with 10 variations each. We categorize errors into perception, decision, action, and no-error cases. For the GPT-4V baseline, we manually assist in action execution, eliminating action errors. This serves as an upper bound for baseline performance. However, even with this enhancement, our RoboEXP consistently shows superior performance.
Figure 6: Experiments on Mobile Robot Scenarios. We visualize the interactive exploration process and the corresponding constructed ACSG in the mobile robot settings. (a) This scenario includes a table, a cabinet, a sofa, multiple small items, and an apple hidden in the top drawer. (b) This scenario involves two chairs, one table, and a reader hidden by the cloth above.
Experiment Settings. To assess our system’s efficacy, we designed five types of experiments, each with 10 different settings varying in object number, type, and layout. We validate the performance of our system in constructing ACSG through quantitative analysis and qualitative demonstrations.
We compare our system with a strong baseline by augmenting GPT-4V with ground truth actions. This baseline operates in a closed-loop fashion, receiving RGB observations from different viewpoints. At each turn, it generates the current scene graph, encompassing hidden objects, and suggests the next action to be taken (refer to the complete prompts in the Appendix). To ensure the baseline is robust, we utilize manual actions as ground truth references for the proposed actions. In contrast, in the exploration experiments described below, all actions from our system are automatically executed by our action module on the physical robot. It is also crucial to note that the output ACSG from our system faithfully aligns with the task requirements. Conversely, for the baseline, we manually construct ACSG based on its actions and the new observations it uncovers. Due to the unstructured nature of the raw scene graph from the baseline, we carefully refine it according to the observable objects, providing an upper-bound performance for comparison during evaluation. (See comparisons with other heuristic-based baselines in Appendix)
We design five key metrics to measure the performance of the interactive exploration task. 1) Success evaluates the success percentage across 10 variants for each task. We define success for each experiment as 1 when the final outputted ACSG exactly matches the GT version, and 0 otherwise; 2) Object Recovery quantifies the percentage of hidden objects successfully identified; 3) State Recovery indicates whether the final state resembles the original state before exploration; 4) Unexplored Space evaluates the percentage of successfully explored need-to-explore space to reduce the robot’s uncertainty about the scene; 5) Graph Edit Distance (GED) measures the disparity between the outputted graph and the GT graph.
Comparison. The quantitative findings presented in Tab. 2 underscore the superior performance of our system compared to the baseline method. Our approach showcases a notable enhancement across all metrics, outperforming the baseline by a considerable margin. The collective assessment of success rate, object recovery, and unexplored space metrics unequivocally validates the efficacy of our system in exploring unfamiliar scenes through interactive processes. It is essential to highlight that, in the case of object recovery, the baseline method may occasionally choose to randomly open certain drawers or doors, revealing objects that have already been discovered. This randomness contributes to a seemingly high object recovery rate but may not necessarily correlate with overall success. The unexplored space metric shows that our system is much more stable in exploring all need-to-explore spaces. Moreover, both the success rate and graph edit distance underscore the close alignment of our system with human actions, highlighting the efficiency of our approach across diverse scenarios. The state recovery metric assesses whether the final state post-exploration resembles the initial state. Our system consistently shows effective state recovery; however, the baseline may trick this metric by opting not to take any action, resulting in an artificially high score in this aspect.
5.2 ACSG for Object Retrieval and Manipulation Tasks
The scenarios depicted in Fig. 1 exemplify the efficacy of our generated output (ACSG) in manipulation tasks. Consider the table-rearranging scenario: without our ACSG, the robot struggles to swiftly prepare the table due to the lack of precise prior knowledge about the location of objects (e.g., the fork stored in the top-left drawer of the wooden cabinet). Beyond comprehensive layout guidance, our ACSG also addresses a crucial question regarding task feasibility for the robot. For instance, if the task requires a spoon but there is no spoon in the scene, the robot recognizes the missing object and asks for human help. In addition to enhancing downstream manipulation tasks, our ACSG possesses the capability to autonomously adapt to environmental changes. In the human intervention setting, our system seamlessly explores newly added components, such as a cabinet, ensuring continuous adaptability. Check our Appendix and supplemental video for more details.
5.3 Extension to Mobile Robot
We also demonstrate the effectiveness of our pipeline using a mobile robot in household scenarios. To adapt to Stretch2, we merely modified our action module to follow the new kinematic structure, while keeping other modules nearly the same. Fig. 6 shows two scenes with hidden objects in drawers and under cloth. Our system is capable of exploring the scene, utilizing manipulation and mobility, to construct the full ACSG and recover the scene. By leveraging the constructed ACSG, we can easily locate the hidden apple and e-reader.
6 Limitations and Future Work
Although our system has proven effective, there is room for improvement. The breakdown of the failure rate in Fig. 4.b suggests that failures primarily arise from detection and segmentation errors within the perception module. To address this issue, we envision two future directions: 1) enhancing the capabilities of visual foundation models for open-world semantic understanding, and 2) utilizing temporal cues and semantic fusion techniques to improve perception robustness through continuous observations. Furthermore, our system would benefit from enhanced LMM capacities and the integration of more sophisticated and generalizable skill modules. Such improvements would enhance both the decision-making and the action execution, thereby further reducing failure cases.
7 Conclusion
We introduced RoboEXP, a foundation-model-driven robotic exploration framework capable of effectively identifying all objects in a complex scene, both directly observable and those revealed through interaction. Central to our system is action-conditioned 3D scene graph, an advanced 3D scene graph that goes beyond traditional models by explicitly modeling interactive relations between objects. Experiments have shown RoboEXP’s superior performance in interactive scene exploration across various challenging scenarios, significantly outperforming a strong GPT4V-based baseline. Notably, the reconstructed action-conditioned 3D scene graph is crucial for guiding complex downstream manipulation tasks, like setting up the table in a mock environment with fridges, cabinets, and drawer sets. Our system and its action-conditioned scene graph lay the groundwork for practical robotic deployment in complex settings, especially in environments like households and offices, facilitating their integration into everyday human activities.
Acknowledgments
Yunzhu Li is partly supported by the Amazon AICE Award. Shenlong Wang is supported by the Amazon AICE Award, IBM IIDAI Grant, Nvidia Hardware Grants, the Insper-Illinois Innovative Grant, the NCSA Faculty Fellow, NSF Award #2331878 and #2312102, and gifts from Intel and Meta. We thank Dorie Wang for generously sharing her toys for our research, Kaifeng Zhang and Yixuan Wang for their kind discussions. This work does not relate to the positions of Shubham Garg and Hooshang Nayyeri at Amazon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
Appendix A Additional Details of Interactive Exploration
Due to space constraints, we did not include the pseudocode of the algorithm proposed in the main paper, but include more details here for clarity. We formulate the interactive scene exploration task into an active perception and exploration problem to construct the action-conditioned 3D scene graph (ACSG).
@@ Appendix A Additional Details of Interactive Exploration
4:      add spatial relations (Algorithm 2) % memory
@@
21:         add action preconditions (Algorithm 3) % memory
@@
In the above algorithm, we have demonstrated how to construct the edges from objects to actions $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{a}}$ and from actions to objects $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{a}}$; however, there is a lack of description for the other two types of edges.
@@
Add Spatial Relations. The logic involves analyzing the spatial relationships among objects using spatial heuristics and incorporating the resulting spatial relation edges between objects $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{o}}$ (Algorithm 2).
@@
Add Action Preconditions. The approach is to assess the feasibility of implementing the actions. We utilize the decision-making module to verify whether there are any prerequisite actions that need to be completed beforehand, and then adjust the plan accordingly (Algorithm 3).
@@
Appendix B Additional Details of RoboEXP System
We provide additional details about our system and each module in it. We then discuss our system’s design for the interactive scene exploration task, focusing on its application in closed-loop exploration processes that may require multi-step or recursive reasoning and handle potential interventions. Additionally, we explain the usage of our proposed ACSG.
B.1 Details of Modules in RoboEXP System
To tackle the interactive exploration task, we present our RoboEXP system, designed to autonomously explore unknown environments by observing and interacting with them. The system comprises four key components: perception, memory, decision-making, and action modules. This closed-loop system ensures the thoroughness of our task in interactive scene exploration.
B.1.1 Perception Module
Raw RGBD images are captured through the wrist camera in different viewpoints and processed by the perception modules to extract scene semantics, including object labels, 2D bounding boxes, segmentations, and semantic features. As mentioned in the main paper, to obtain per-instance CLIP features, we implement a strategy similar to the one proposed by Jatavallabhula et al. [94]. Specifically, we extend the local-global image feature merging approach by incorporating additional label text features to augment the semantic CLIP feature for each instance. Furthermore, we exclusively focus on instance-level features, disregarding pixel-level features, thereby accelerating the entire semantic feature extraction process.
B.1.2 Memory Module
The memory module is designed to construct our ACSG of the environment by assimilating observations over time. For the low-level memory, to ensure stable instance merging from 2D to 3D, we employ a similar instance merging strategy as presented in Lu et al. [95], consolidating observations from diverse RGBD sources across various viewpoints and time steps. In contrast to the original algorithm, which considers only 3D IoU and semantic feature similarity we additionally incorporate label similarity and instance confidence. To enhance algorithm efficiency, we represent low-level memory using a voxel-based representation with filtering designs, which allows for more efficient computation and cleaner memory updates. Meanwhile, given the crowded nature of objects in our tabletop setting, we have implemented voxel-based filtering designs to obtain a cleaner and more complete representation of the objects for storage in our memory.